Consumption =445+0,75y investment =250 1.1 calculate the value of autonomous spending 1.2calculate the value of multiplier 1.3 calculate equilibrium level of income
1
Expert's answer
2021-10-25T09:50:41-0400

"C=A+MD"

Where,

C- consumer spending.

A- autonomous consumption.

M- Marginal Prospensity to Consume.

D- Disposable Income.


We are given:

"Consumption=445+0.75y"

Therefore:

(1)

Value of autonomous spending=445.

(2)

Value of multiplier = marginal propensity to consume = 0.75.

(3)

Equilibrium Level of Income:

"Y= Consumption + Investment"

"Y=445+0.75Y+250"

"Y-0.75Y=445+250"

"0.25Y=695"

"Y= 2,780."